OpenStudy (anonymous):

solve by the elimination process 2x + 3y = 4 4x + 6y = 8 what is the solution of the system

OpenStudy (anonymous):

hi jazil Do you have a selection of answer? i think its inconsistent here 2x + 3y = 4 if you want to eliminate x here, multiply by 2 4x + 6y =8 (2x + 3y = 4)2 4x + 6y = 8 4x + 6y = 8 4x + 6y = 8 ------------ 0 +0 = 0

OpenStudy (anonymous):

so in other words there is no solution or there are definitely many solutions which one

OpenStudy (anonymous):

it means that the two equations are the equally the same..either you substitute x or y, it will come up to be equal

OpenStudy (anonymous):

can you graph this two equations?

OpenStudy (anonymous):

graph them and you will find them to be the same equation

OpenStudy (anonymous):

it does not tell you to graph I have to tell what the solution of the system is by selecting a. the solution is b. there are infinitely many solutions c. there is no solution

OpenStudy (anonymous):

ok c, no solution

OpenStudy (anonymous):

ok c, no solution

OpenStudy (anonymous):

ok thats what i picked can you help me with this one please solve and graph 1 - 3x >or=to 10 and 7x-5 > 30

OpenStudy (anonymous):

1 - 3x >or=to 10 and 7x-5 > 30 - 3x >or=to 9 and 7x > 35 - x >or=to 3 and x>5 x <or=to -3

OpenStudy (anonymous):

did you get it jazil?

OpenStudy (anonymous):

Would the correct ans be {x|x-3<=tox<5} or {x|x<=to-3} those are the two choices out of four to choose from and if one is correct where would the graph be meaning what end is the bracket and arrow

OpenStudy (anonymous):

can you give me the other two choices pls?

OpenStudy (anonymous):

sure one is O with a line though it and {x|x>5}

OpenStudy (anonymous):

{x|x<=to-3} {x|x>5} these two are correct answer as we calculated there....lol

OpenStudy (anonymous):

understand how i did the problem there and youl get them all.....lol

OpenStudy (anonymous):

yep i was checking that out very well done so on the graph i would choose the one with two parts correct with an arrow pointing to the left with a backet on right end and parenthese on left for other and arrow pointing to the right on end of this one am i right

OpenStudy (anonymous):

yes correct put the arrow to the left for x <or=to -3 and the other x>5 it must be > than 5

OpenStudy (anonymous):

and the other x>5 it must be > than 5, to the right

OpenStudy (anonymous):

thanks thats what i did

OpenStudy (anonymous):

ok jazil good luck and have fun.....

OpenStudy (anonymous):

thanks for all your help tonight i learn a lot from you in such a short while i've been on for a little bit getting tired so i say nite nite

OpenStudy (anonymous):

ok ywc, good nite,,,lol...